OpenStudy (anonymous):

a silver ball weighing 50g is removed from a furnace and dropped into a 350g of water at 0°C. if the equilibrium temperature is 10°C and the specific heat capacity of platinum is 230J〖kg〗^(-1)〖°C〗^(-1), what is the temperature of the furnace? neglect the effect of the mass of the calorimeter.

OpenStudy (anonymous):

315

OpenStudy (anonymous):

can you explain? i really have no idea how to do this. :(

OpenStudy (anonymous):

q=mct so,heat absorbed by water=350*1*10=3500 and this(3500) is heat released by silver ball so, 3500=0.23*50*t t=314 this t is change in temp so, temp of ball=314=10=315

OpenStudy (anonymous):

oops.......t=304 so temp=304+10=314

OpenStudy (anonymous):

thank you very much! :) god bless you :)

OpenStudy (shane_b):

Well, that could not be more wrong :)